### v1

# Uncertainty-aware data assimilation through variational inference

Abstract
Data assimilation, consisting in the combination of a dynamical model with a set of noisy and incomplete observations in order to infer the state of a system over time, involves uncertainty in most settings. Building upon an existing deterministic machine learning approach, we propose a variational inference-based extension in which the predicted state follows a multivariate Gaussian distribution. Using the chaotic Lorenz-96 dynamics as a testing ground, we show that our new model enables to obtain nearly perfectly calibrated predictions, and can be integrated in a wider variational data assimilation pipeline in order to achieve greater benefit from increasing lengths of data assimilation windows. Our code is available at https://github.com/anthony-frion/Stochastic_CODA.
Index Terms—  Data assimilation, uncertainty quantification, variational inference, deep learning
1 Introduction
In many geoscience problems, one has access to an accurate description of a dynamical system $\mathcal{M}$ of a studied state $\mathbf{x}_{t}\in\mathbb{R}^{n}$, but the capacity for directly measuring the full state is lacking. Instead, one can rely on data assimilation, in which partial information obtained from observations $\mathbf{y}_{t}$ enable the identification of $\mathbf{x}_{t}$ using the state-space equations:
$\mathbf{x}_{t+1}=\mathcal{M}(\mathbf{x}_{t})+\boldsymbol{\eta}_{t},$ (1)
$\mathbf{y}_{t}=\mathcal{H}_{t}(\mathbf{x}_{t})+\boldsymbol{\epsilon}_{t}.$ (2)
The observation operator $\mathcal{H}_{t}$ represents a functional relationship from $\mathbf{x}_{t}$ to $\mathbf{y}_{t}$ which in the general case cannot be inverted since $\mathbf{y}_{t}$ is typically much lower dimensional than $\mathbf{x}_{t}$. $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\epsilon}_{t}$ respectively denote model and observation errors. They are often assumed to follow centered Gaussian distributions that do not depend on time $t$, and are thus described by their covariance matrices $\boldsymbol{\Sigma_{\eta}}$ and $\boldsymbol{\Sigma_{\epsilon}}$.
In this context, we seek to estimate the joint posterior probability distribution of the state $\mathbf{x}_{1:T}$ given a set of observations $\mathbf{y}_{1:T}$ covering the same time span, i.e. $p_{\mathbf{x}}(\mathbf{x}_{1:T}|\mathbf{y}_{1:T})$. This data assimilation task is an inverse problem [1], where the state equation (1) is used (alongside a prior distribution of $\mathbf{x}_{1}$) to build a prior on $p_{\mathbf{x}}(\mathbf{x}_{1:T})$. Popular classes of methods for solving it include Kalman smoothers [2] and ensembles thereof [3], as well as variational methods such as 4D-Var [4]. Many recent papers have proposed machine learning-based approaches (e.g. [5, 6, 7, 8, 9, 10]), often with deterministic outputs (i.e. estimating only the maximum a posteriori of $p_{\mathbf{x}}(\mathbf{x}_{1:T}|\mathbf{y}_{1:T})$). While most of these methods rely on the supervised training of a neural network using ground truth state trajectories $\mathbf{x}_{1:T}$ or analysis performed by a classical assimilation algorithm, we leverage a recently proposed unsupervised learning method [11] which trains a model directly from noisy and incomplete observations $\mathbf{y}_{1:T}$ without ever requiring access to the corresponding states $\mathbf{x}_{1:T}$.
2 Variational CODA
The model $G_{\theta}$ from [11], named Combined Optimization of Dynamics and Assimilation (CODA), takes as input a window of observations $\mathbf{y}_{t-w:t+w}$ and returns an estimate $\hat{\mathbf{x}}_{t}=G_{\theta}(\mathbf{y}_{t-w:t+w})$ of the most likely state at the center of the window given these observations. While in some cases one could perform a supervised training of $G_{\theta}$ using the mean squared error between its predictions $\hat{\mathbf{x}}_{t}$ and the true state $\mathbf{x}_{t}$, the authors instead use the following unsupervised training loss:
$L(\theta)=\mathbb{E}_{t}\bigr[\sum_{i=0}^{h}||\mathbf{y}_{t+i}-\mathcal{H}_{t+i}\circ\mathcal{M}^{(i)}(\hat{\mathbf{x}}_{t})||^{2}\\ +\lambda||\hat{\mathbf{x}}_{t+h}-\mathcal{M}^{(h)}(\hat{\mathbf{x}}_{t})||^{2}\bigr],$ (3)
where $\circ$ denotes composition of functions, and $h$ is a hyperparameter representing the horizon of the prediction. The first term is an observation error, measuring the agreement of the advancement of the predicted $\hat{\mathbf{x}}_{t}$ by $i$ time steps with the corresponding observations $\mathbf{y}_{t+i}$, for $i$ between 0 and $h$. The second term is a regularization promoting the self-consistency of $G_{\theta}$ by comparing its time-propagated prediction to a prediction made by itself with a similar window of observations $h$ time steps later. The hyperparameter $\lambda$ is fixed and enables adjustment of the relative weight of these two loss terms.
Here, we modify $G_{\theta}$ so that, instead of returning a pointwise estimate $\hat{\mathbf{x}}_{t}$ of the state, it returns the parameters $\boldsymbol{\mu}_{t}\in\mathbb{R}^{n},\boldsymbol{\sigma}_{t}\in\mathbb{R}^{n}$ of a diagonal Gaussian distribution, through
$G_{\theta}(\mathbf{y}_{t-w:t+w})=(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t}).$ (4)
One possibility to train this new variational model would be to introduce a prior on $\boldsymbol{{\mathbf{x}}}_{t}$ and leverage a Kullback-Leibler divergence loss with respect to this prior, as explained in [12]. We opt instead for an adaptation of the self-consistency loss from equation (3) using the negative log-likelihood of samples from $\hat{\mathbf{x}}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})$ [^1] From here on, $\boldsymbol{\Sigma}_{t}$ is a diagonal matrix with diagonal coefficients $\boldsymbol{\sigma}_{t}$. when propagated $h$ steps forward, with regards to the distribution $\mathcal{N}(\boldsymbol{\mu}_{t+h},\boldsymbol{\Sigma}_{t+h})$ inferred at time $t+h$. Thus, equation (3) becomes
$L(\theta)=\mathbb{E}_{t,\hat{\mathbf{x}}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})}\bigr[\sum_{i=0}^{h}||\mathbf{y}_{t+i}-\mathcal{H}_{t+i}\circ\mathcal{M}^{(i)}(\hat{\mathbf{x}}_{t})||^{2}\\ -\lambda\;\log p(\mathcal{M}^{(h)}(\hat{\mathbf{x}}_{t})|\boldsymbol{\mu}_{t+h},\boldsymbol{\Sigma}_{t+h})\bigr].$ (5)
It should be noted that this choice is a natural extension of equation (3), as the mean squared error can be understood as a negative log-likelihood with an identity covariance matrix. Besides, without the self-consistency term, the loss function would tend to encourage the predicted variances to tend to 0, actually leading to a deterministic model in practice. Thus, the choice of $\lambda$ in equation (5) appears to be critical for the calibration of the uncertainty of our probabilistic model.
3 Training a stochastic CODA model
Following the experiments of [11], we work on the Lorenz-96 dynamical system [13], a popular benchmark in data assimilation, which models the evolution of a meteorological quantity on a latitude circle (i.e. the domain is periodic). It is composed of $n$ variables $x_{1},...,x_{n}$, each evolving as
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F.$ (6)
We use $n=40$ variables and a forcing parameter $F=8$, resulting in chaotic dynamics with a doubling time of approximately $0.42$ time units (i.e. 2.1 days). The system is numerically integrated with a time step of $\delta t=0.01$ (i.e. 1.2 hours) using the Runge-Kutta 4 integration scheme. The observation operator $\mathcal{H}$ randomly masks 75% of the variables at each time step. The remaining variables are observed with a standard Gaussian observation error $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0}_{n},\mathbf{I}_{n})$.
Our primary evaluation metric is the continuous ranked probability score (CRPS) [14], averaged over time and over the $n$ dimensions of the system. The CRPS is a proper scoring rule enabling to compare the cumulative distribution function $F$ of a (one-dimensional) predicted probability distribution with a pointwise ground truth state $x_{*}$ as follows:
$\text{CRPS}(F,x_{*})=\int_{-\infty}^{\infty}[F(x)-\mathds{1}_{x\geq x_{*}}]^{2}\text{d}x.$ (7)
The CRPS is a generalization of the mean absolute error to stochastic predictions. Indeed, for an ensemble $(x_{1},...,x_{M})$ of $M$ equiprobable member predictions, we have:
$\text{CRPS}=\frac{1}{M}\sum_{i=1}^{M}|x_{*}-x_{i}|-\frac{1}{2}\frac{1}{M^{2}}\sum_{i=1}^{M}\sum_{j=1}^{M}|x_{i}-x_{j}|,$ (8)
which trivially reduces to the mean absolute error when $M=1$. Equation (8) can be used in practice to estimate the CRPS of any probability distribution by sampling from it.
As secondary metrics, we use the spread and skill of the predictions. The spread of a predicted distribution is its standard deviation, and its skill is the root mean squared error of its mean with respect to the ground truth state. The spread and skill are expected to match on average when an accurate and well-calibrated posterior distribution has been computed, and one can test for this with the spread-skill plot. This consists in binning the spread values (obtained with varying inputs) in a histogram and computing the associated skills for each of these bins. Two synthetic metrics that one can derive from it are the spread-skill ratio (SSRAT) and spread-skill reliability (SSREL). The SSRAT is defined as the global ratio between the spread and the skill, and it has an ideal value of $1$, with values above indicating underconfidence and values below indicating overconfidence. The SSREL is the sum of absolute differences between the binned spread and skill values. Thus, it depends on the binning process and has an ideal value of 0. The interested reader can refer to [15] for more extensive descriptions of the CRPS, SSRAT and SSREL.
We train stochastic models on three datasets, respectively built from trajectories integrated over $10^{4}$, $3\times 10^{5}$ and $3\times 10^{6}$ time steps. Therefore, we hereafter refer to them as the small, medium and big datasets.
The benchmarked models are:
Variational. The method from section 2.
Dropout. An adaptation of the deterministic CODA model from [11], with the simple addition of dropout [16] on its output layer, randomly masking out neurons with a probability $p$ that is adjusted to minimize the CRPS and obtain calibrated uncertainties. Following [17], we use dropout during both training and inference, which results in a stochastic behavior mimicking Bayesian neural networks [18].
Ensembling. We independently train 5 dropout models, which differ only by their random initializations, as suggested by [19], and combine their predictions when testing.
Dataset size Small Medium Big
Variational CRPS 0.295 $\mathbf{0.195}$ $\mathbf{0.168}$
SSRAT 0.731 $\mathbf{1.035}$ $\mathbf{1.000}$
SSREL 0.105{0.105} $\mathbf{0.012}$ $\mathbf{0.010}$
Dropout CRPS 0.283{0.283} 0.206 0.187
SSRAT 0.780{0.780} 1.040 0.970
SSREL 0.103{0.103} 0.074 0.067
Ensembling CRPS $\mathbf{0.246}$ 0.197 0.176
SSRAT $\mathbf{1.075}$ 1.075 1.093
SSREL $\mathbf{0.064}$ 0.073 0.062
Table 1: Performance of three stochastic prediction methods, on three different dataset sizes. For CRPS and SSREL, lower values are better. For SSRAT, values closer to 1 are better.
The results of these 3 methods are summarized in table 1, and the spread-skill plots of their instances trained on the big dataset are shown on figure 1. From these results, one can first see that the CRPS of all 3 benchmarked methods strongly decreases as the amount of training data increases. It further appears that ensembling 5 dropout models consistently reduces CRPS relative to a single dropout model. As can be seen from the fact that the spread-skill plot of the variational method is close to the 1:1 line in figure 1, this method obtains a significantly better SSREL than the other two methods when the training data is abundant. One can also note that ensembling dropout models improves the skill of the predictions but has a low impact on the spread, hence the increased SSRAT. While the SSREL of the variational method (and, to a lesser extent, of a single dropout model) clearly improves with larger amounts of training data, the same cannot be said for the ensemble of dropout models, for which there is no clear tendency. Overall, the variational method performs best except when trained on the smallest dataset. This can be explained by the ability of dropout layers to prevent overfitting in neural networks [16], which is indeed most important when training on low amounts of data. From these results, it seems likely that an ensemble of variational CODA models would perform even better than a single model, especially when re-tuning $\lambda$ in equation (4) to obtain a well calibrated ensemble.
Fig. 1: Spread-skill plots for the variational, dropout and ensembling models, trained on the big dataset. The inset spread frequencies plot indicates the relative weights of the dots of the main plot when computing the SSREL. The 1:1 line represents a perfect calibration.
4 Using a trained model for 4D-Var
In the previous section, we have trained a stochastic CODA model that produces nearly perfectly calibrated diagonal Gaussian estimates of the state of a system. This model, once trained, is much more computationally efficient than classical data assimilation methods. However, it computes an estimate of a state at time $t$ using a relatively small window of observations (e.g. 65 time steps for our model trained on the biggest dataset) and it does not directly simulate the known dynamics of the system at inference time. This means that one can expect improved performance with a costlier method, leveraging observations over much longer assimilation windows. Thus, in this section, we integrate a pre-trained instance of our stochastic CODA model into a classical 4D-Var computation scheme for assimilation on long observation windows. We show that this approach enables to improve the performance of deterministic data assimilation compared to a direct use of CODA, and more importantly compared to a similar 4D-Var scheme that does not leverage the CODA outputs.
Let us denote by $\mathbf{y}_{-w:T+w}$ a window of observations from the Lorenz-96 system, following the same characteristics as in section 3. We use a pre-trained variational CODA model to get initial predictions $(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ on the corresponding state variables through equation (4), for every time t from 0 to $T$. Then, in its more general form, the assimilation cost that we aim to minimize in our 4D-Var scheme is expressed as
[…]
where $\mathbf{x}||^{2}_{\mathbf{A}}=\mathbf{x}^{\intercal}\mathbf{A}^{-1}\mathbf{x}$ is a weighted Euclidean norm. $\alpha,\beta,\gamma$ are hyperparameters used to adjust the relative weights of the 4 terms of the cost. While $\alpha$ can take any positive value, we will only consider values of 0 or $1$ for $\beta$ and $\gamma$. When $\beta=1$ and $\gamma=0$, equation (9) can be recognized as an instance of the classical weak-constraint 4D-Var cost (see e.g. [20]). It allows computing the maximum a posteriori of the distribution $p(\mathbf{x}_{0:T}|\mathbf{y}_{0:T})$ when the background prior distribution of the initial state is a Gaussian $\mathbf{x}_{0}\sim\mathcal{N}(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$, the model error in equation (1) is $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0}_{n},\alpha^{-1}\mathbf{I}_{n})$ and the observation error follows a standard Gaussian distribution. While the background prior on the initial state for each assimilation window is usually derived from analysis of a previous window, using stochastic CODA as a prior makes use of the first $w$ observations in the present window as well. Thus, these observations are used twice—once when setting up the prior, and again when minimizing the cost in equation (9). Besides, when additionally setting $\gamma=1$, we make use of a “foreground prior”, which we introduce analogously to the background prior in order to guide the assimilation cost at the end of the observation window. Note that using a foreground prior is not standard in variational data assimilation.
Fig. 2: Mean squared error obtained by our 4 variants of 4D-Var, as a function of the assimilation window length. For each length, we use 10 different windows, and report the mean plus/minus standard deviation of the errors.
We test 4 variants of the assimilation procedure, on 7 different lengths $T$ of the observation window. For each value of $T$, we use 10 different sets of states and observations, and report the mean and standard deviation of the mean squared errors of each variant over these inverse problem instances.
The values of $T$ that we consider are [$1000$, $2000$, $5000$, $10^{4}$, $2\cdot 10^{4}$, $5\cdot 10^{4}$, $10^{5}$]. We use $\alpha=10^{7}$ in all experiments: a very high value since the dynamical model is in fact perfectly known here. The assimilation cost is minimized using automatic differentiation, and the optimizer is limited-memory BFGS, run for 5000 iterations with the default parameters from Pytorch. The 4 tested variants are:
Nearest init. A minimization of the cost from equation (9) with $\beta=\gamma=0$ and an initial value of $\mathbf{x}_{0:T}$ obtained by, for each variable and each time step, copying the available observation that is the closest in time for this variable. Like the other variants, this initialization makes use of a few observations outside of the assimilation window. This heuristic initialization was selected as a baseline as it surprisingly performed better than slightly more complex approaches such as linear interpolation and Cressman interpolation [21].
CODA init. A variant that similarly sets $\beta=\gamma=0$ but uses the mean predictions $\boldsymbol{\mu}_{t}$ of the pre-trained CODA model to initialize $\mathbf{x}_{0:T}$ in equation (9).
CODA init and background prior. Same as CODA init, but additionally using $\beta=1$, so that the predicted mean and variance of the initial state are explicitly used to define the background prior term in the assimilation cost.
CODA init, background and foreground prior. Same as the previous variant, but additionally using $\gamma=1$.
In figure 2, we show the results obtained by these different variations. One can first observe that all of them strongly benefit from increasing assimilation window lengths. CODA init performs better than Nearest init in all cases, with a gap strongly increasing with the window length. CODA init and background prior enables significant improvements, especially for the shorter window lengths. Finally, the inclusion of a foreground prior results in a marginal additional reduction of the mean squared error for shorter window sizes. Thus, both the background and foreground priors are most useful when the amount of observed data is relatively low.
Fig. 3: Visualization of a slice of an assimilation window with $T=10^{5}$ time steps. First row: ground truth state trajectory. Second row: Noisy and incomplete observations used in the 4D-Var cost. Third row: resulting predicted trajectory. Fourth row: differences between the predicted and true states.
In figure 3, we show some results over a slice of one of our longest assimilation windows ($T=10^{7}$), assimilated with the CODA init, foreground and background prior approach. One can see that the ground truth is very well reconstructed using only sparse and noisy observations.
5 Conclusion
In this paper, we showed that it was possible to obtain nearly perfectly calibrated uncertainties in data assimilation with an unsupervised training of neural network for variational inference. We also demonstrated how such a pre-trained model can help improve the reconstruction ability of a classical weak-constraint 4D-Var method. However, in this experiment, we leverage stochastic estimates of the state in order to ultimately obtain deterministic predictions, and thus designing revised 4D-Var-like methods with stochastic outputs is a natural extension of this work. Furthermore, while we have only considered data assimilation with a perfectly known dynamical model, the CODA framework can also address more general tasks where the dynamics are partly unknown, and uncertainty-aware resolution of these tasks would certainly be of interest. Finally, it remains to be seen how our methods would perform on systems of larger scale than Lorenz-96.